6. The small OAB lever, which is placed in the pressure cooker, works in accordance with the moment rule. The force F of the vapor pressure creates the opposite moment in relation to the axis of rotation O, of the lever. Calculate this force, when it is known that the mass of sphere B is m = 200 g, while OA = 0.5 cm and OB = 5 cm.


1
Expert's answer
2021-03-04T15:19:16-0500
"F_1r_{OB}=F_2r_{OA},""F_2=\dfrac{F_1r_{OB}}{r_{OA}}=\dfrac{0.2\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot5\cdot10^{-2}\ m}{0.5\cdot10^{-2}\ m}=19.6\ N."